OpenStudy (anonymous):

We know our customer adds 18% to cost. So if cost is $200 then 18% added would equal $236. How do we get back to cost when we only know the $236 (ie back out the 18% that was added)? We can't figure out the math. We tried mulitply $236 *.82 (the inverse but that doesn't work).

OpenStudy (anonymous):

x=200, with added value of 18% you get 200*(1.18) = 236. However, you now have y=236, you know that 18% was added to some value to obtain 236. so x*(1.18) = 236. So what it looks like is ( x + .18*x) = 236 --> x*(1.18) = 236. so x = 236/1.18

OpenStudy (anonymous):

You are the best and I'm a dummy. Thanks for making me look good at work!